OpenStudy (anonymous):

How do I solve: |2x + 5 | - 9 = 12

OpenStudy (anonymous):

|2x + 5 | - 9 = 12 |2x+5| = 12+9 |2x+5| = 21 since it is an absolute value you have to have two answers for x 2x+5 = 21 2x = 21-5 2x = 16 x = 8 AND 2x+5 = -21 2x = -26 x = -26/2 x = -13

terenzreignz (terenzreignz):

too bad this is no longer my favourite kind of absolute value equation... oh well... make it so that the absolute value part is alone on one side of the equation.

OpenStudy (johnweldon1993):

you're going to have 2 answers....add 9 to both sides...then you'll have 2x + 5 = 21 and 2x + 5 = -21

OpenStudy (johnweldon1993):

@terenzreignz what's your favorite now? lol

terenzreignz (terenzreignz):

This http://openstudy.com/study#/updates/51430325e4b04cdfc581822d

OpenStudy (johnweldon1993):

lol ahh so ones that confuse everyone haha....i remember when i first looked at one of those and just had the blankest exression until i realized...hmm..it's impossible lol

terenzreignz (terenzreignz):

Confuses for a bit, but then again, if all your questions were like that, you'd do the test in less than minute... :D